An 4.7-kg object moving at 12.2 m/s catches up and collides with an 22.3-kg object moving at 2.1 m/s in the same direction. After the collision the 4.7-kg object is observed to be moving in the opposite direction with the speed of 2.6 m/s.

How much kinetic energy was lost in this system?
Answer #1

since no external force so momentum will be conserved

4.7*12.2 + 22.3*2.1 = -4.7*2.6 + 22.3 *x

x= 5.21928251121 m/s

so after collision 22.3 Kg object will move with v =5.21928251121 m/s

KE loss = 0.5 * 22.3*(5.21928251121)^2 + 0.5* 4.7*(2.6)^2 -(0.5*22.5*(2.1)^2 + 0.5 * 4.7 * (12.2)^2)

= 79.7643542602  J